Consider a 6×10−2 M  solution of the weak base methylamine.

The pKb  of methylamine is 3.34  

You can search for the structure of this compound online, although precise knowledge of the structure is not needed.

Calculate the pH and the concentration of all species present in equilibrium.

1.Calculate the concentration of   methylammonium ions.

2.Calculate the concentration of   methylamine in equilibrium.
Answer #1

Methylamine is CH3NH2

In aqueous solution it dissociates as

CH3NH2 + H2O CH3NH3+ + OH-

ICE table is

X is degree of dissociation

[CH3NH2] [CH3NH3+] [OH-]
Initial C 0 0
Change - Cx + Cx +Cx
Equilibrium C-Cx Cx Cx

Then, Kb = [CH3NH3+] [OH-]/[CH3NH2]

= [Cx*Cx/(C-Cx)]

Kb = Cx2/ (1-x)

As methyl amine is weak base , 1>>x

Kb = Cx2

Given , pKb = 3.34

Or, - logKb = 3.34

Then, Kb = 0.000457

Or, Cx2 = 0.000457

Or, X2 = (0.000457/C) = (0.000457/6*10-2) = 0.0076

Or, x = 0.087

Then, pOH = - log [OH-] = - log[Cx] = - log(6*10-2*0.087)

= -log(0.00522) = 2.282

Then pH = 14 - pOH = 14 - 2.282 = 11.717

Concentration of methylammonium ions is [CH3NH3+] = [OH-] = 0.00522 M

and, concentration of methylamine in equilibrium is

[CH3NH2] = (C - Cx) = (6*10-2 - 0.00522) = 0.05478 M